OpenStudy (anonymous):

What is the reduce or simplified form of these radicals: 1.)sqrt of 12 + stqrt of 108 3.)cube-root of 16 * the cube-root 3

OpenStudy (raden):

sqrt(12) = sqrt(4*3) = sqrt(4) * sqrt(3) = 2 sqrt(3) while sqrt(108) = sqrt(36*3) = sqrt(36) * sqrt(3) = 6 sqrt(3) add them, it will be 2sqrt(3) + 6sqrt(3) = 8 sqrt(3)

OpenStudy (anonymous):